# Severity Prediction in Mental Health: LLM-based Creation, Analysis, Evaluation of a Novel Multilingual Dataset

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly being integrated into various medical fields, including mental health support systems. However, there is a gap in research regarding the effectiveness of LLMs in non-English mental health support applications. To address this problem, we present a novel multilingual adaptation of widely-used mental health datasets, translated from English into six languages (Greek, Turkish, French, Portuguese, German, and Finnish). This dataset enables a comprehensive evaluation of LLM performance in detecting mental health conditions and assessing their severity across multiple languages. By experimenting with GPT and Llama, we observe considerable variability in performance across languages, despite being evaluated on the same translated dataset. This inconsistency underscores the complexities inherent in multilingual mental health support, where language-specific nuances and mental health data coverage can effect the accuracy of the models. Through comprehensive error analysis, we emphasize the risks of relying exclusively on large language models (LLMs) in medical settings (e.g., their potential to contribute to misdiagnoses). Moreover, our proposed approach offers significant cost savings for multilingual tasks, presenting a major advantage for broad-scale implementation.
Severity Prediction in Mental Health: LLM-based Creation, Analysis, Evaluation of a Novel Multilingual Dataset
Konstantinos Skianis1, John Pavlopoulos2, A. Seza Doğruöz3 1University of Ioannina, Greece 2Athens University of Economics and Business, Greece 3Universiteit Gent, Belgium Corresponding author: kskianis@cse.uoi.gr
@@ 1 Introduction
Large language models (LLMs) (e.g., GPT-3, BERT, and their multilingual variants) are trained on vast amounts of text data and demonstrate an impressive ability to generate human-like text, perform complex language tasks Devlin et al. (2019); Brown et al. (2020). One of the most promising applications of LLMs is in the healthcare domain, where they have the potential to assist medical specialists with the detection, diagnosis, and treatment of health conditions.
Mental health disorders such as depression, anxiety, and post-traumatic stress disorder (PTSD) are widespread across the globe, presenting major public health challenges World Health Organization (2021). Traditional diagnostic approaches typically involve self-reported questionnaires, clinical interviews, and standardized assessments performed by trained professionals Kessler and Üstün (2004). However, these processes are also time-consuming and labor intensive which leads to delays in treatment and increased waiting times for the patients.
Therefore, leveraging large language models (LLMs) to detect mental health symptoms from textual data presents a promising alternative. These models have the ability to process vast amounts of textual content (e.g., social media posts, online forum discussions, personal narratives), and detect specific linguistic indicators related to mental health issuesGuntuku et al. (2019); Chancellor et al. (2019). In this way, health professionals can be assisted with the early detection and intervention of mental health symptoms, reduce delays for patients whose symptoms might otherwise go unnoticed.
However, most studies about applying LLMs into mental health research are still in English and there is a need to explore this research area for other languages Raihan et al. (2024).
Our study fills these gaps with the following research questions:
@@
  • We create a novel multilingual dataset (covering English, Turkish, French, Portuguese, German, Greek, and Finnish) derived from user-generated social media content, which is automatically translated from English using an LLM. [^1] The dataset is publicly available at the following link: https://github.com/y3nk0/multilingual-mental-severity-prediction.
@@
To the best of our knowledge, this is the first study to utilize LLMs for creating and evaluating multilingual adaptations of existing mental health NLP datasets, specifically designed to detect the severity of mental health conditions. Our study pushes forward the creation of AI-powered tools and datasets for mental health diagnosis that are both inclusive and adaptable, emphasizing the critical need for approaches that account for cultural and linguistic diversity in mental health care, especially in non-English contexts with cost-efficient methods.
@@ 2 Related work
LLMs excel at detecting mental health symptoms with greater precision by leveraging their advanced understanding of context and semantics. Notable examples include BioBERT Lee et al. (2020) and ClinicalBERT Huang et al. (2019), which have been pre-trained on biomedical literature and clinical notes. Additionally, models like MentalBERT Ji et al. (2022), DisorBERT Aragon et al. (2023), SuicidalBERT Anonymous , MentalLLM Xu et al. (2024) and MentalLlama Yang et al. (2023) are specifically trained on mental health-related social media data. Nevertheless, using LLMs for detecting mental health issues may introduce a number of advantages, as well as disadvantages De Choudhury et al. (2023). Additionally exploiting social network posts represents a promising frontier in both artificial intelligence and mental health care Harrigian et al. (2020). LLMs (e.g., GPT-4 and BERT) can analyze vast amounts of textual data from social media platforms automatically to identify linguistic patterns associated with mental health conditions such as depression, anxiety, and suicidal ideation Ji et al. (2023).
@@ 3 Proposed Methodology
We present a comprehensive methodology for leveraging LLMs and social media posts (in English), in order to translate them into multiple languages and predict mental health conditions accordingly.
@@
Dep-Severity Naseem et al. (2022) Depression 4 3553 Minimum (2587)
@@
Sui-Twi Mirza Ibtihaj et al. (2020) Suicide 3 1300 Not Suicidal (515)
Suicidal Ideation (455)
Potential Suicide Attempt (330)
@@ 4.1 Datasets
In order to test the proposed methodology we are using the following datasets:
@@
Sui-Twi
This dataset (Mirza Ibtihaj et al., 2020) contains texts from Twitter (in English), including 1,300 suicidal ideation tweets. The authors gathered posts that contained the word ‘suicide’ and labeled them manually into three classes ”not suicidal", ”indicative of suicidal ideation", and ”indicative of a potential suicide attempt". In this dataset, the classes are balanced in terms of frequency of occurence.
@@
Dep-Severity “Categorize the following text with 1 of the 4 depression severity levels (0: Minimum, 1: Mild, 2: Moderate, 3: Severe)"
Sui-Twi “Categorize to 0 if the text is not suicidal, to 1 if the text is almost related to suicidal ideation, and to 2 if the text indicates a potentially suicidal user"
@@ 4.2 Settings
not suicidal 0.78 0.53 0.63 0.58 0.44 0.50 0.58 0.66 0.62 0.63 0.53 0.58 0.56 0.67 0.61 0.58 0.67 0.63 0.57 0.52 0.55
ideation 0.46 0.78 0.58 0.36 0.70 0.47 0.43 0.40 0.42 0.43 0.57 0.49 0.39 0.42 0.41 0.49 0.50 0.49 0.35 0.42 0.38
potential 0.69 0.36 0.47 0.54 0.04 0.07 0.54 0.47 0.50 0.54 0.41 0.47 0.51 0.31 0.39 0.65 0.47 0.55 0.49 0.39 0.43
Macro avg 0.64 0.56 0.56 0.49 0.39 0.35 0.51 0.51 0.51 0.53 0.51 0.51 0.49 0.47 0.47 0.57 0.55 0.55 0.47 0.45 0.45
Table 4: GPT-3.5 with 0-shot learning on Sui-Twi, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best macro-averaged F1 across languages is underlined.
@@
not suicidal 0.62 0.88 0.73 0.83 0.61 0.70 0.67 0.80 0.73 0.66 0.71 0.68 0.62 0.87 0.72 0.64 0.81 0.71 0.62 0.84 0.71
ideation 0.48 0.32 0.38 0.46 0.60 0.52 0.51 0.49 0.50 0.43 0.55 0.49 0.42 0.38 0.40 0.56 0.44 0.49 0.41 0.32 0.36
potential 0.75 0.61 0.67 0.45 0.45 0.45 0.78 0.60 0.68 0.79 0.41 0.54 0.78 0.38 0.52 0.81 0.73 0.77 0.69 0.51 0.58
Macro avg 0.62 0.60 0.59 0.58 0.55 0.56 0.66 0.63 0.64 0.63 0.55 0.57 0.61 0.54 0.55 0.67 0.66 0.66 0.57 0.56 0.55
Table 6: GPT-4o-mini with 0-shot learning on Sui-Twi, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
@@
English French Portuguese German
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
not suicidal 0.76 0.54 0.63 0.44 0.50 0.47 0.67 0.58 0.62 0.49 0.63 0.55
ideation 0.43 0.40 0.41 0.38 0.15 0.21 0.47 0.50 0.49 0.47 0.25 0.33
potential 0.45 0.68 0.54 0.23 0.37 0.28 0.48 0.53 0.50 0.39 0.45 0.42
Macro avg 0.55 0.54 0.53 0.35 0.34 0.32 0.54 0.54 0.54 0.45 0.45 0.43
Table 8: Llama3.1 with 0-shot learning on Sui-Twi, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
@@
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
not suicidal 0.72 0.11 0.20 0.55 0.26 0.35 0.47 0.03 0.05 0.59 0.21 0.31 0.50 0.16 0.25 0.72 0.03 0.07 0.39 0.09 0.14
ideation 0.37 0.81 0.51 0.36 0.25 0.29 0.68 0.09 0.16 0.37 0.36 0.37 0.51 0.33 0.40 0.17 0.00 0.00 0.34 0.10 0.15
potential 0.39 0.26 0.31 0.36 0.80 0.50 0.27 0.98 0.42 0.30 0.61 0.40 0.32 0.82 0.46 0.26 0.99 0.41 0.27 0.87 0.42
Macro avg 0.49 0.40 0.34 0.42 0.44 0.38 0.47 0.37 0.21 0.42 0.39 0.36 0.44 0.44 0.37 0.38 0.34 0.16 0.33 0.35 0.24
Table 10: GPT-3.5 with 1-shot learning on Sui-Twi by measuring Precision, Recall, and F1 per class per language. The last row shows macro averaged scores. The best F1 per class is shown in bold and the best average F1 across languages is underlined.
@@ 4.3 Preliminary Prompting
  • Observable shifts in behavior or routines, such as disrupted sleep or changes in appetite.
@@
    Not Suicidal:
    • General Tone: Neutral, positive, or typical negative expressions not related to self-harm or life-threatening situations.
    • Content: Discussions about daily activities, hobbies, interests, work, or social interactions without any reference to self-harm or death.
    • Language: Absence of words or phrases related to self-harm, death, hopelessness, or helplessness.
@@
    Indicative of Suicidal Ideation:
    • Expression of Hopelessness: Phrases indicating a lack of hope for the future, such as “I can’t see any way out", “Things will never get better", or "There’s no point in trying".
    • Self-Hatred or Worthlessness: Statements like “I’m a burden", “Everyone would be better off without me," or "I hate myself".
    • Mentions of Death or Dying: Expressions about wanting to die or not wanting to live, such as “I wish I were dead", “I want to end it all", or “Life isn’t worth living".
    • Social Withdrawal: Indications of isolating oneself from others, such as “I don’t want to see anyone", or “I prefer to be alone all the time".
@@
    Indicative of a Potential Suicide Attempt:
    • Detailed Plans or Intentions: Descriptions of specific methods or plans to commit suicide, such as “I have pills ready," “I’m going to jump off a bridge," or “I’ve written my goodbye letter."
    • Goodbye Messages: Posts that read like final farewells, such as “Goodbye everyone, it’s been a pleasure," or “I’m sorry for everything, I won’t be around much longer."
    • Previous Attempts: References to past suicide attempts, particularly if combined with current suicidal thoughts, such as “I’ve tried before and failed, but not this time."
    • Lethal Means: Mentions of acquiring or possessing means to commit suicide, like “I bought a gun," “I have enough pills," or “I found a rope."
When asked, the LLM presents a detailed description of the three classes, which seems promising.
@@ 4.4 Experimental results
Initially, we experimented with the data in their source language (English), to set the baseline performance. That is, no translation step has been performed at this stage. When using GPT-3.5, the first column in Tables 3 and 4 presents the results on depression and suicide detection tasks respectively. The best F1 in both tasks is achieved for the lowest severity/indication. In Dep-Severity, the best performance is achieved for the two edges of lowest (F1=0.48) and highest severity (F1=0.26). In Sui-Twi, on the other hand, F1 drops as the suicidal likelihood increases, from 0.73 to 0.67. The task of Dep-Severity is more challenging (F1=0.17) compared to Sui-Twi (F1=0.56). When using the superior GPT-4o-mini (Tables 6 and 5), F1 increases to 0.27 and to 0.59, respectively. This can be attributed to the fact that detecting specific levels of depression is harder than in the case of suicidal ideation, as depression severity levels are considered less distinct. Therefore, it is probably more challenging for the LLM to distinguish.
@@
GPT-3.5 has a varying performance when operating on content translated from English to target languages. In Dep-Severity (Table 3), performance drops in Turkish (-6), Portuguese (-1), Greek (-1), and Finish (-2). On the other hand, the performance increases in French (+3) and German (+7). In Sui-Twi (Table 3), performance drops in all target languages, with the drop for Turkish (-17) being the most considerable one. This can be explained by the limited number of Turkish resources in health, especially in the domain of mental health Çöltekin et al. (2023).
GPT-4o-mini in Dep-Severity (Table 5) performs equally well in Portuguese and German (as in English), but this is not true for the rest of the target languages. In Sui-Twi, performance drops for most target languages. However, we observe again the phenomenon of better performance for a target language, namely for French (+5) and Greek (+7), as was the case for GPT-3.5 in French and German. Therefore, we find that by translating to French, performance increases for both models. The best performance is found for Greek (F1=0.66) and the worst for German and Finnish (-4).
@@
Figures 2 and 3 present F1 in boxplots, summarising the performance of GPT-3.5 across languages per task. In Dep-Severity, the lowest (minimum) severity class is characterised by a wide range, from 0.15 to 0.25. The whiskers are also wide, from 0.05 (Turkish) to 0.38 (German). Mild severity is mishandled across languages while for the classes of moderate and severe, the model achieves a relatively high score across languages.
@@
Figure 3: F1 of GPT-3.5 across languages (vertically) per class (box) for Sui-Twi.
In Sui-Twi, the best performance is achieved for the “non-suicidal" category. This is also the major class while slightly worse performance is achieved for the other two classes. For all three classes the performance is similarly distributed across languages.
@@
Tables 9 and  10 present the results when 1-shot learning is employed. In English, performance drops considerably in Sui-Twi and remains similarly low in Dep-Severity. Again, performance is better for a target language, compared to the performance for the source language, but considerably lower (or equal) than what achieved with zero-shot learning. Better performance could be observed if more examples were provided, but this would also increase the cost of the experiments by approx. doubling the size of the prompt.
@@
Tables 7 and 8 show the performance of Llama-3.1 in both tasks, using 0-shot, for the languages it was available for. Although it is better in English compared to GPT-3.5 (both 0- and 1-shot), it is slightly worse in target languages. We observe high variability in Sui-Twi, with French being the worst and Portuguese the best, but not in Dep-Severity. In both tasks, the best performance is not achieved in English.
@@
Figure 4: An example of translation (English to Greek) from the suicide dataset. The original words and their translations are colored similarly.
Lastly, Figure 4 presents two instances of the suicide dataset. Regarding the first example, while the translation from English to Greek appears to be correct in terms of conveying the same message, the English text is classified as "Potential suicide attempt", while the Greek text is classified as "Suicidal ideation". The LLM might be interpreting the context differently due to nuances in the languages. In English, "attempted suicide" explicitly suggests a recent suicide attempt, which clearly falls under "Potential suicide attempt." In Greek, the model might be weighing other parts of the text more heavily, leading to a classification of "Suicidal ideation." This could be due to differences in the training data, language model intricacies, or contextual emphasis.
For the second example, we observe that the translation is accurate and conveys the same message as the original English text. Both the English and Greek texts are classified consistently and correctly as "suicidal ideation".

### v2

# Building Multilingual Datasets for Predicting Mental Health Severity through LLMs: Prospects and Challenges

@@ Abstract
Large Language Models (LLMs) are increasingly being integrated into various medical fields, including mental health support systems. However, there is a gap in research regarding the effectiveness of LLMs in non-English mental health support applications. To address this problem, we present a novel multilingual adaptation of widely-used mental health datasets, translated from English into six languages (e.g., Greek, Turkish, French, Portuguese, German, and Finnish). This dataset enables a comprehensive evaluation of LLM performance in detecting mental health conditions and assessing their severity across multiple languages. By experimenting with GPT and Llama, we observe considerable variability in performance across languages, despite being evaluated on the same translated dataset. This inconsistency underscores the complexities inherent in multilingual mental health support, where language-specific nuances and mental health data coverage can affect the accuracy of the models. Through comprehensive error analysis, we emphasize the risks of relying exclusively on LLMs in medical settings (e.g., their potential to contribute to misdiagnoses). Moreover, our proposed approach offers significant cost savings for multilingual tasks, presenting a major advantage for broad-scale implementation.
Building Multilingual Datasets for Predicting Mental Health Severity through LLMs: Prospects and Challenges
Konstantinos Skianis, A. Seza Doğruöz#, John Pavlopoulos Department of Computer Science and Engineering, University of Ioannina, Greece # LT3, IDLab, Universiteit Gent, Belgium Department of Informatics, Athens University of Economics and Business, Greece Archimedes/AthenaRC, Greece kskianis@cse.uoi.gr  as.dogruoz@ugent.be  annis@aueb.gr
@@ 1 Introduction
Warning: this paper contains text examples that are negative, depressive, or adverse.
Large language models (LLMs) (e.g., GPT-3, BERT, and their multilingual variants) are trained on vast amounts of text data and demonstrate an impressive ability to generate human-like text, perform complex language tasks Devlin et al. (2019); Brown et al. (2020). One of the most promising applications of LLMs is in the healthcare domain, where they have the potential to assist medical specialists with the detection, diagnosis, and treatment of health conditions of patients.
Mental health disorders (e.g., depression, anxiety, and post-traumatic stress disorder (PTSD)) are widespread across the globe, presenting major public health challenges World Health Organization (2021). Traditional diagnostic approaches typically involve (manual)-reported questionnaires, clinical interviews, and standardized assessments performed by trained professionals Kessler and Üstün (2004). However, these processes are also time-consuming and labor intensive, leading to delays in treatment and increased waiting times for the patients.
Therefore, leveraging large language models (LLMs) to detect mental health symptoms from textual data has presented a promising alternative. These models, for instance, have the ability to process vast amounts of textual content (e.g., social media posts, online forum discussions, personal narratives), and detect specific linguistic indicators related to mental health issues Guntuku et al. (2019); Chancellor et al. (2019). In this way, health professionals can be assisted with the early detection and intervention of mental health symptoms, reduce delays for patients whose symptoms might otherwise go unnoticed.
However, most studies about applying LLMs to mental health research are still in English and there is a need to explore this research area for other languages Raihan et al. (2024). Our study fills these gaps with the following research questions:
@@
  • We create a novel multilingual dataset (covering English, Turkish, French, Portuguese, German, Greek, and Finnish) derived from user-generated social media content, which is automatically translated from English using an LLM. [^1] Code and some results are uploaded as supplementary material. The dataset will be made publicly available upon acceptance.
@@
To the best of our knowledge, this is the first study to utilize LLMs for creating and evaluating multilingual adaptations of existing mental health NLP datasets, specifically designed to detect the severity of mental health conditions. Our study pushes forward the creation of inclusive and adaptable AI-powered tools and datasets for mental health diagnosis, emphasizing the critical need for approaches that account for cultural and linguistic diversity in mental health care, especially in non-English contexts with cost-efficient methods.
@@ 2 Related work
LLMs excel at detecting mental health symptoms with greater precision by leveraging their advanced understanding of context and semantics. Notable examples include BioBERT Lee et al. (2020) and ClinicalBERT Huang et al. (2019), which have been pre-trained on biomedical literature and clinical notes. Additionally, models like MentalBERT Ji et al. (2022), DisorBERT Aragon et al. (2023), SuicidalBERT Anonymous , MentalLLM Xu et al. (2024) and MentalLlama Yang et al. (2023) are specifically trained on mental health-related social media data. Nevertheless, using LLMs for detecting mental health issues may introduce a number of advantages, as well as disadvantages De Choudhury et al. (2023). Additionally, exploiting social network posts represents a promising frontier in both artificial intelligence and mental health care Harrigian et al. (2020). LLMs (e.g., GPT-4 and BERT) can analyze vast amounts of textual data from social media platforms automatically to identify linguistic patterns associated with mental health conditions such as depression, anxiety, and suicidal ideation Ji et al. (2023).
@@ 3 Proposed Methodology
We present a comprehensive methodology for leveraging LLMs and social media posts (in English), to translate them into multiple languages and predict mental health conditions accordingly.
@@
Dep-Sev Naseem et al. (2022) Depression 4 3553 Minimum (2587)
@@
C-ssrs Gaur et al. (2019) Suicide 5 500 Supportive (108)
Indicator (99)
Ideation (171)
Behavior (77)
Attempt (45)
@@ 4.1 Datasets
In order to test the proposed methodology (above), we use the following datasets:
@@
C-SSRS
The C-SSRS (Columbia-Suicide Severity Rating Scale) dataset, created by Gaur et al. (2019), is a collection of structured data used to assess suicidal ideation and behavior. It is based on the Columbia-Suicide Severity Rating Scale, a widely used tool in clinical and research settings to evaluate the severity and frequency of suicidal thoughts, past attempts, and related risk factors. The dataset includes several variables (e.g., suicidal ideation intensity, past attempts, preparatory behaviors, and protective factors) allowing researchers and mental health professionals to identify high-risk individuals.
@@
Dep-Sev “Categorize the following text with 1 of the 4 depression severity levels (0: Minimum, 1: Mild, 2: Moderate, 3: Severe)”
C-ssrs “Categorize the content into the following classes based on the user’s mental state: 1. Supportive, 2. Indicator, 3. Ideation, 4. Behavior, 5. Attempt.”
@@ 4.2 Settings
Supportive 0.27 0.13 0.18 0.40 0.02 0.04 0.30 0.19 0.24 0.36 0.04 0.07 0.38 0.27 0.31 0.39 0.35 0.37 0.29 0.09 0.14
Indicator 0.21 0.76 0.33 0.20 0.95 0.34 0.21 0.72 0.32 0.21 0.92 0.34 0.27 0.57 0.37 0.22 0.49 0.31 0.22 0.44 0.30
Ideation 0.45 0.26 0.33 0.25 0.05 0.08 0.40 0.21 0.28 0.30 0.09 0.14 0.40 0.49 0.44 0.39 0.42 0.40 0.36 0.56 0.44
Behavior 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.50 0.03 0.05 0.00 0.00 0.00 0.00 0.00 0.00
Attempt 0.00 0.00 0.00 0.25 0.02 0.04 0.00 0.00 0.00 1.00 0.02 0.04 1.00 0.04 0.09 0.00 0.00 0.00 0.00 0.00 0.00
Macro avg 0.19 0.23 0.17 0.22 0.21 0.10 0.18 0.22 0.17 0.37 0.21 0.12 0.51 0.28 0.25 0.20 0.25 0.22 0.17 0.22 0.18
Table 4: GPT-3.5-turbo with 0-shot learning on C-ssrs, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
@@
Supportive 0.40 0.65 0.50 0.40 0.69 0.51 0.43 0.60 0.50 0.49 0.57 0.53 0.44 0.56 0.50 0.45 0.54 0.49 0.39 0.70 0.50
Indicator 0.26 0.52 0.34 0.27 0.36 0.31 0.27 0.44 0.33 0.29 0.55 0.38 0.24 0.46 0.32 0.27 0.53 0.36 0.21 0.29 0.25
Ideation 0.48 0.20 0.29 0.46 0.35 0.40 0.46 0.32 0.37 0.43 0.26 0.33 0.42 0.36 0.39 0.46 0.46 0.46 0.44 0.37 0.40
Behavior 1.00 0.01 0.03 0.31 0.06 0.11 0.67 0.03 0.05 0.50 0.01 0.03 0.00 0.00 0.00 0.00 0.00 0.00 0.50 0.01 0.03
Attempt 0.39 0.47 0.42 0.30 0.22 0.26 0.31 0.44 0.37 0.34 0.62 0.44 0.42 0.22 0.29 0.57 0.09 0.15 0.35 0.13 0.19
Macro avg 0.51 0.37 0.32 0.35 0.34 0.32 0.43 0.37 0.33 0.41 0.40 0.34 0.30 0.32 0.30 0.35 0.32 0.29 0.38 0.30 0.27
Table 6: GPT-4o-mini with 0-shot learning on C-ssrs, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
@@ 4.3 Preliminary Prompting
  • Observable shifts in behavior or routines (e.g., disrupted sleep or changes in appetite).
@@
    Supportive (Encouragement and Positive Reinforcement). Some key linguistic features that may appear include expressions of empathy, encouragement, positivity and use of supportive phrases (e.g., ”You are not alone”, ”Stay strong”, ”I’m here for you”, ”You matter”).
@@
    Indication (Signs of Distress or Risk). Linguistic features may include the use of worrying expressions (e.g., ”I feel lost”, ”I don’t know what to do”, ”Nothing makes sense anymore”, ”I’m so tired”; indirect mentions of mental health struggles: ”I’m always exhausted”, ”Everything is meaningless”, ”I just want it to stop”).
@@
    Ideation (Expressions of Suicidal Thoughts). Key linguistic features may include direct expressions of suicidal thoughts (e.g., ”I wish I didn’t exist”, ”I want to end it all”, ”I don’t see the point of living”; questions about death and suicide: ”What happens after we die?”, ”Is there peace after death?”; use of hopelessness in an extreme form: ”Nothing will ever change”, ”I just want to disappear”).
  4. 4.
    Behavior (Actions Related to Suicide). Key linguistic features include descriptions of planning or preparation (e.g., ”I’ve been researching painless ways to go”, ”I wrote my letter today”, ”This might be my last post”, ”I have the pills ready”; mentions of giving away possessions or saying goodbye: ”I just want to thank everyone for being in my life”).
  5. 5.
    Attempt (Reports of Actual Suicide Attempts). Key linguistic features include explicit mentions of self-harm or suicide attempts (e.g., ”I tried last night but failed”, ”I was in the hospital after overdosing”, ”I attempted but someone stopped me”, ”I survived but I don’t know if I should be happy about it”) and descriptions of medical intervention (e.g., ”They pumped my stomach”, ”They found me just in time”, ”I’m in the ER right now”).
@@ 4.4 Experimental results
Initially, we experimented with the data in their source language (English), to set the baseline performance. In other words, no translation step has been performed at this stage. When using GPT-3.5, the first column in Tables 3 and 4 presents the results on depression and suicide detection tasks respectively. The best F1 in both tasks is achieved for the lowest severity/indication (minimum depression and supportive for suicide). In Dep-Severity, the best performance is achieved for the two edges of lowest (F1=0.48) and highest severity (F1=0.26). In C-ssrs, on the other hand, F1 drops from 0.18 to 0.00. When using the superior GPT-4o-mini (Tables 5 and 6 ), F1 increases to 0.27 and to 0.32, respectively. Detecting specific levels of suicidal ideation is harder than in the case of depression, as they may be considered less distinct. Therefore, it could be probably more challenging for the LLM to distinguish between them, also shown by the results for some classes.
@@
GPT-3.5 has a varying performance when operating on content translated from English to target languages. In Dep-Severity (Table 3), performance drops in Turkish (-6), Portuguese (-1), Greek (-1), and Finish (-2). On the other hand, the performance increases in French (+3) and German (+7). In C-ssrs (Table 4), performance drops in all target languages, with the drop for Turkish (-7) being the most considerable one. This can be explained by the limited number of Turkish resources in health, especially in the domain of mental health Çöltekin et al. (2023).
GPT-4o-mini in Dep-Severity (Table 5) performs equally well in Portuguese and German (as in English). However, this is not true for the rest of the target languages. In C-ssrs, performance remains similar for most target languages. However, we observe again the phenomenon of better performance for a target language, namely for French (+1) and Portuguese (+2), as was the case for GPT-3.5 in German. In general, GPT-4o-mini is much more stable and consistent for correct prediction.
@@
Figures 2 and 3 present F1 in boxplots, summarising the performance of GPT-3.5 across languages per task. In Dep-Severity, the lowest (minimum) severity class is characterized by a wide range, from 0.15 to 0.25. The whiskers are also wide, from 0.05 (Turkish) to 0.38 (German). Mild severity is mishandled across languages while for the classes of moderate and severe, the model achieves a relatively high score across languages.
@@
Figure 3: F1 of GPT-4o-mini across languages (vertically) per class (box) for C-ssrs.
In C-ssrs, the best performance is achieved for the “Supportive” category. This is also the second major class, while slightly worse performance is achieved for the other classes. For all classes, apart from “Attempt” the performance is similarly distributed across languages.
@@
Table 8 presents the results when 1-shot learning is employed. In English, performance remains similarly low in Dep-Severity. Again, performance is better for a target language (German), compared to the performance for the source language. In general, we observe an increase going from 0-shot to 1-shot. Better performance could be observed if more examples were provided. However, this would also increase the cost of the experiments by approx. doubling the size of the prompt.
@@
Table 7 shows the performance of Llama-3.1 in Dep-Severity using 0-shot, for the languages it was available for. Although the model performs better in English compared to GPT-3.5 (both 0- and 1-shot), it is slightly worse in target languages. Notably, the best performance is not achieved in English.
Back-translation
To assess the effect of translation, we translated the already translated texts back to English and repeated the experiments for GPT-3.5-turbo with 0-shot on C-ssrs. Table 9 shows that the prediction increased when translating back to English. The highest gain (from 0.17; see Table 4) is observed when translating to Portuguese or Finnish and back (+0.05) while +0.03 is reported for Turkish and +0.02 for the rest. Hence, we observe that translating before predicting assists the prediction performance of the LLM. We present some representative examples in the Appendix.
Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
Supportive 0.43 0.33 0.38 0.37 0.26 0.30 0.44 0.34 0.39 0.37 0.31 0.34 0.36 0.19 0.24 0.40 0.34 0.37
Indicator 0.22 0.77 0.35 0.22 0.75 0.34 0.23 0.76 0.36 0.20 0.64 0.31 0.22 0.79 0.34 0.23 0.74 0.35
Ideation 0.43 0.19 0.27 0.42 0.20 0.27 0.40 0.22 0.28 0.41 0.22 0.29 0.45 0.22 0.29 0.41 0.21 0.28
Behavior 1.00 0.01 0.03 0.67 0.03 0.05 0.00 0.00 0.00 1.00 0.01 0.03 0.50 0.01 0.03 0.00 0.00 0.00
Attempt 0.00 0.00 0.00 0.00 0.00 0.00 1.00 0.04 0.09 0.00 0.00 0.00 0.50 0.02 0.04 1.00 0.04 0.09
Macro avg 0.42 0.26 0.20 0.33 0.25 0.19 0.42 0.27 0.22 0.40 0.24 0.19 0.40 0.24 0.19 0.41 0.27 0.22
Table 9: Back-translation and prediction with GPT-3.5-turbo with 0-shot learning on C-ssrs, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
@@
Figure 4: An example of translation (English to Greek) from the C-ssrs dataset with GPT-4o-mini.
Lastly, Figure 4 presents two instances of the suicide dataset. Regarding the first example, while the translation from English to Greek appears to be correct in terms of conveying the same message, the English text is classified as ”Supportive”, while the Greek text is classified incorrectly as ”Indicator”. In this case, GPT could interpret the context differently due to the nuances in the languages. In Greek, the model might be weighing other parts of the text more heavily, leading to a classification of ”Indicator”. This could be due to differences in the training data, the intricacies of the language model, or contextual emphasis.
For the second example, we observe that the translation is accurate and conveys the same message as the original English text. Nevertheless, the English text was incorrectly classified as level 2 (indicator). However, the Greek translation was correctly classified as level 3 (ideation). Here, we see that detecting different levels is inherently difficult. The original English post could be mispredicted as it is quite small and generic.